OpenStudy (anonymous):

are THESE fractions equivalent? 84/112 and 168/192

OpenStudy (mathteacher1729):

Do you know how to factor a number like 84? (write it as the product of prime numbers)

OpenStudy (anonymous):

84/112 = 3/4 168/192 = 7/8 since 3/4 =/= 7/8 they aren't equivalent